# Rajatsubhra Chakraborty Arkaprava Sinha  Dominick Reilly  Manish Kumar Govind Pu WangFrancois BremondSrijan Das UNC Charlotte  Inria  Université Côte d’Azur

### v1

@@ Abstract
Large Language Vision Models (LLVMs) have demonstrated effectiveness in processing internet videos, yet they struggle with the visually perplexing dynamics present in Activities of Daily Living (ADL) due to limited pertinent datasets and models tailored to relevant cues. To this end, we propose a framework for curating ADL multiview datasets to fine-tune LLVMs, resulting in the creation of ADL-X, comprising 100K RGB video-instruction pairs, language descriptions, 3D skeletons, and action-conditioned object trajectories. We introduce LLAVIDAL, an LLVM capable of incorporating 3D poses and relevant object trajectories to understand the intricate spatiotemporal relationships within ADLs. Furthermore, we present a novel benchmark, ADLMCQ, for quantifying LLVM effectiveness in ADL scenarios. When trained on ADL-X, LLAVIDAL consistently achieves state-of-the-art performance across all ADL evaluation metrics. Qualitative analysis reveals LLAVIDAL’s temporal reasoning capabilities in understanding ADL. The link to the dataset is provided at: https://adl-x.github.io/
Figure 1: Comparison of LLVM vs LLAVIDAL : In real world scenarios, web-video trained models struggle to understand Activities of Daily Living due to the subtle nuances in the video, whereas our ADL-X trained LLAVIDAL model triumphs in understanding complex human-object interactions.
@@ 1 Introduction
Human cognitive perception integrates information from multiple sensory modalities to form a unified representation of the world [1]. Towards emulating human cognitive perception in digital intelligence, initial efforts focused on integrating vision and language modalities [2, 3, 4, 5, 6]. Subsequently, the success of LLMs like GPT [7], PALM [8], BLOOM [9] led to the introduction of multimodal conversational models[10, 11, 12, 13, 14, 15, 16] that combine image pixels and LLMs, we dub as Large Language-Vision Language Models (LLVMs). However, these image-based LLVMs lack the capability for complex reasoning and interactions, particularly in understanding spatio-temporal relationships involved in human activities. In this study, we investigate the understanding of Activities of Daily Living (ADL) videos by LLVMs, which present various challenges including multiple exo-centric viewpoints, fine-grained activities with subtle motion, complex human-object interactions, and long-term temporal relationships. We envision that LLVMs capable of addressing these challenges will significantly influence the future intelligent systems, particularly in healthcare applications such as eldercare monitoring, cognitive decline assessment, and robotic assistance development.
@@
Recently,  [17, 18, 19, 20, 21, 22, 23] have integrated videos into LLMs, leading to the development of video-based LLVMs capable of capturing spatio-temporal features. However, these models are predominantly trained on large-scale web videos [24, 25, 26, 27, 28], which mainly consists of sports clips, movie excerpts, and instructional videos. These videos, typically filmed by professionals, follow strict temporal sequences in closely controlled background (e.g., Paragliding). The evident temporal structure and scene semantics in such videos facilitate spatial understanding within LLVMs, as shown in 1. In contrast, ADL videos pose additional challenges, characterized by temporal unstructuredness where diverse actions may unfold concurrently within a single sequence [29]. For instance, a person cooking could intermittently engage in unrelated activities like making a phone call or drinking water, disrupting the linear progression of the composite action cooking. Consequently, existing LLVMs trained on web videos struggle to capture such visually perplexing dynamics inherent in ADL scenarios. Moreover, unlike specialized video architectures designed for understanding ADL [30, 31, 32, 33, 34, 35, 36], these LLVMs lack explicit utilization of cues like 3D poses or object encodings, which are crucial for understanding ADL. These cues aid in learning view-invariant representations and capturing fine-grained details essential for interpreting complex human activities. Hence, the current limitations in understanding ADL stem from the lack of instruction tuning of LLVMs on real-world multiview ADL datasets captured in indoor settings and the simplistic design of LLVMs with holistic operations.
To this end, we propose a framework of curating ADL videos for instruction tuning LLVMs. This framework introduces the ADL-X dataset, comprising 100K untrimmed RGB video-instruction pairs, 3D poses (P), language descriptions, and action-conditioned object trajectories (see Table 1). We then introduce the Large LAnguage VIsion model for Daily Activities of Living (LLAVIDAL), trained on ADL-X, which integrates videos, 3D poses, and object cues into the LLM embedding space. Our study explores various strategies for integrating 3D pose information and human-object interactions within LLVMs, demonstrating that language contextualized features extracted from 3D poses and object trajectories can effectively be integrated into LLAVIDAL. Furthermore, we introduce a benchmark ADL Multiple Choices Question (ADLMCQ), specifically designed to evaluate the effectiveness of LLVMs for ADL. ADLMCQ includes action recognition (ADLMCQ-AR) and action forecasting (ADLMCQ-AF), assessed through a multiple choice question-answering task. We also evaluate existing LLVMs for generating video description of ADL scenes and compare their performance with LLAVIDAL. Our empirical findings indicate that LLAVIDAL with object cues, outperforms other LLVMs, including those trained on datasets of ten times the size, on the ADL benchmarks.
To summarize our contributions:
@@
  • We introduce ADL-X, the first multiview RGBD instruction ADL dataset, curated through a novel semi-automated framework for training LLVMs.
  • LLAVIDAL is introduced as the first LLVM tailored for ADL, incorporating 3D poses and object cues into the embedding space of the LLM.
  • A new benchmark, ADLMCQ, is proposed for an objective evaluation of LLVMs on ADL tasks, featuring MCQ tasks for action recognition & forecasting.
  • Exhaustive experiments are conducted to determine the optimal strategy for integrating poses or objects into LLAVIDAL. Evaluation of existing LLVMs on ADLMCQ and video description tasks reveals that LLAVIDAL trained on ADL-X significantly outperforms baseline LLVMs.
Table 1: Video Instruction Dataset Comparison.
Dataset Modalities Subjects Multiple Videos QA Pairs Atomic Actions Temporal Object Type
Views per Vid Rand. Traj.
TimeIT[21] RGB+L NA No 173000 173K Medium No No Web
VideoChat[17] RGB+L NA No 8196 11K Low No No Web
Valley[26] RGB+L NA No 64,687 65K Low No No Web
VideoChatGPT [20] RGB+L NA No 27,801 100K Medium No No Web
ADL-X RGB+P+L 106 Yes 16,343 100K High Yes Yes ADL
2 Semi-automated Framework for generating ADL Video-instructions Pairs
This section describes the data curation framework employed for the creation of a novel dataset, ADL-X. This dataset specifically caters to the instruction tuning of LLVMs within the ADL domain. ADL-X comprises video recordings of ADLs. To enrich the dataset and facilitate LLM training, question-answer (QA) pairs were generated from a corpus of long-form ADL videos. These QA pairs target various aspects of the ADLs, including: human pose configuration, objects relevant to the human actions, scene appearance, and the fine-grained actions performed. We hypothesize that incorporating such instructional tuning during the LLVM training process will promote alignment of visual tokens within the LLM’s embedding space. ADL-X represents a comprehensive ADL dataset encompassing various modalities: - RGB videos, 3D poses, Language descriptions, object tracklets. This rich dataset offers a valuable tool for evaluating the capabilities of LLVMs in tasks related to ADLs, including description, recognition, and anticipation.
A critical characteristic of ADL videos lies in the inherent spontaneity of the actions performed. Unlike scripted scenarios [25, 37, 38], fine-grained actions within ADLs often occur randomly. To capture this essential characteristic within our dataset, we curated ADL-X from NTU RGB+D 120 dataset [39]. This selection was motivated by the dataset’s focus on ADL videos and its inherent diversity in terms of actions, subjects, and camera viewpoints. Also, this data curation framework could be extended to any existing trimmed/untrimmed ADL datasets [40, 41, 42]. Below, we elaborate the steps involved in building the ADL-X in a chronological order.
Person-centric Cropping. ADL tasks necessitate a focus on the individual performing the actions, the actions themselves, and the human-object interactions. To achieve this targeted focus within the data curation framework, we implemented a person-centric cropping strategy leveraging the pose information captured through Kinect sensors [43]. By using the pose information in each frame of the NTU RGB+D 120 dataset, we are able to detect and crop out the person(s) performing the actions. This cropping process effectively reduces the amount of background information present in the videos, eliminating data irrelevant to the target ADLs. This step is crucial as existing ADL datasets often contain extensive background information that is not relevant to the actions being performed. The presence of such extraneous information can significantly hinder subsequent stages within the data curation framework.
Figure 2: Dataset Curation Pipeline: We employ CogVLM[44] as our person-centric image captioner and GPT 3.5 Turbo[7] as our summarizer and QA generator.
Stitching shorts clips. To capture the inherent randomness of real-world ADLs, we constructed a set of 160 composite action sequences. These sequences were generated by prompting a GPT to combine individual actions from the original NTU RGB+D 120 dataset’s list of 120 actions (denoted as $A_{1}$, $A_{2}$, …, $A_{120}$). An example sequence structure could be represented as $A_{1}\rightarrow A_{3}\rightarrow A_{17}$. Following these generated composite action sequences, we temporally stitched together short video clips ($clip_{j}^{a}$, where $a$ is the action class) from the NTU dataset. This stitching process ensured that all clips within a video belonged to the same subject and camera view, maintaining coherence in the resulting video sequence. For instance, a stitched video sequence might be represented as [$clip_{r1}^{1}\hskip 7.22743ptclip_{r2}^{3}\hskip 7.22743ptclip_{r3}^{17}$] where $r1$, $r2$, $r3$ represent unique clip identifiers within the dataset for the specific subject performing the actions (actions 1, 3, and 17, respectively). The intentional randomness of the generated action sequences reflects the unstructured flow of actions encountered in ADL. To further enhance diversity and ensure no bias towards specific subject-action combinations, we shuffled both the action sequences and the subject assignments. This process resulted in the creation of 16,343 stitched videos with an average 5 actions per video.
Frame Level Captioning and Dense Descriptions. This step is the process of generating weak pseudo-labels for automated instruction tuning of the LLVM with the curated dataset. An image captioning model CogVLM [44] is employed to automatically generate frame-level captions for the stitched ADL videos at a rate of $0.5fps$. These captions are subsequently compiled into a dictionary linking each frame identifier to its corresponding description. To enhance the reliability of the pseudo-labels, we implemented an action-conditioned filtering while generating the video descriptions. The dictionary with the frame descriptions, along with the action labels present in the stitched videos, are then used to prompt a GPT 3.5 turbo model to generate a cohesive structured description of the entire stitched video, constrained to a maximum of 300 words. This step leverages the known action labels associated with each video to remove irrelevant noise potentially introduced during the caption generation process. We evaluated various image captioning models, including BLIP-2 [45], and InstructBLIP [46] for frame-level caption generation. However, CogVLM is ultimately chosen due to its ability to generate denser and appropriate descriptions. Please refer to Appendix H for our detailed prompting strategy in generating the descriptions.
Generating QA Pairs. LLVMs necessitate training data in the form of question-answer (QA) pairs. To generate domain-specific QA pairs for ADL, we leverage the dense video descriptions obtained in the previous step as illustrated in Figure 2. An instruction template (detailed in Appendix H) guides GPT-3.5 in formulating questions across various categories relevant to ADL. These categories include: video summary, performed actions, spatial details, human-object interactions and other video-specific inquiries. Through this prompting approach, we curate a dataset of 100K video instruction pairs, namely ADL-X, for the stitched ADL videos. These QA pairs benefit from the detailed descriptions and person-centric cropping, resulting in reduced LLM hallucinations compared to other existing methods [17, 20].
Notably, the framework employed for constructing ADL-X from trimmed, labeled action videos can be generalized to other existing datasets. This generalization paves the way for efficient training of domain-specific LLVMs.
Figure 3: Overview of LLAVIDAL, which utilizes an LLM to integrate multiple modalities, including video, pose, and object features. Videos are represented by embeddings obtained from a VLM, poses are processed through (PoseLM), and object embeddings are obtained through (ObjectLM). These embeddings are projected into the LLM space, where they are concatenated with tokenized text queries for instruction tuning.
3 LLAVIDAL: An LLVM for ADL
LLAVIDAL is a large language vision model designed to align ADL videos with an LLM to generate meaningful conversation about the daily activities performed by humans. This model, similar to Video-ChatGPT [20] and LLaVA [18], integrates a visual encoder with the Vicuna language decoder [47] and is fine-tuned on instructional language-vision data. Unlike Video-ChatGPT [20] and LLaVA [18], LLAVIDAL leverages the random temporal structure present in ADL-X and incorporates additional data modalities such as 3D human poses and human-object interaction cues. This allows LLAVIDAL to generate accurate conversations that are not only contextually appropriate but also temporally aligned with the human activities depicted in the input video. This section will first present a background of LLVM models to align videos with LLMs. Then, we will outline the strategies employed to integrate 3D poses and object interaction cues within the language space of the LLM for enhanced understanding of videos featuring ADL. Subsequently, we will describe the training architecture of LLAVIDAL.
3.1 Background: LLVM
@@
Following [20], given an input video denoted by $\nu_{i}\in\mathbb{R}^{T\times H\times W\times C}$, where $T$ represents the frames encoded using a pretrained vision-language model (VLM) CLIP-L/14 [2] to obtain frame-level embeddings for the video, $x_{i}\in\mathbb{R}^{T\times h\times w\times D}$, with $D$ as the embedding dimension, and $h=H/p$, $w=W/p$ representing the dimensions adjusted by patch size $p$. Temporal and spatial features are extracted by aggregating these frame-level embeddings along the respective dimensions. The video-level features, $V_{i}\in\mathbb{R}^{F_{\nu}\times D_{\nu}}$, are obtained by concatenating the temporal and spatial features, where $F_{v}$ represents the spatio-temporal tokens and $D_{\nu}$ is the video feature dimension. The video features are projected into the LLM embedding space using a linear projection layer $\mathcal{T}_{v}$. Thus, we obtain input tokens $Q_{v}$ for the video features:
$Q_{v}=\mathcal{T}_{v}(V_{i})\in\mathbb{R}^{F_{v}\times K}$ (1)
@@
3.2 3D Poses for LLAVIDAL
ADL are rich in actions that primarily involve the movements of critical body parts or joints. The dataset ADL-X includes 3D human poses, which can be utilized to incorporate human kinematics and view-invariant features into the input embedding space of a LLM. These poses can be integrated into the LLM input space in several ways: as an additional text query $Q_{t}$ for instruction tuning of the LLM, by deriving language descriptions of joint movements to provide context for the LLM, or through features extracted using a suitable pose-language encoder.
Poses as QA. We input the 3D joint coordinates alongside the associated human action from the video into GPT-3.5 Turbo [7], which generates a general description of the pose. This description is then re-fed into GPT-3.5 Turbo to generate two QA pairs that provide detailed explanations of the action’s motions. These QA pairs are subsequently added to the set of text queries $Q_{t}$ in our training set for instruction tuning the LLM.
Poses as Context. To extract contextual information from human poses, we initially identify five peripheral joints — the head, right hand, left hand, right knee, and left knee — due to their significant contribution to motion in various actions. Using GPT-3.5 Turbo, we generate descriptions of the motion for each of these joints based on their trajectories throughout the video, specifically focusing on how the coordinates of these five joints evolve. The generated descriptions, denoted as $Q^{p}_{t}$, are subsequently appended to the text query $Q_{t}$, incorporates these pose descriptions as additional contextual information. This enriched query $Q_{t}^{new}=[Q^{p}_{t}\hskip 3.61371ptQ_{t}]$ is then employed for instruction tuning of the LLAVIDAL.
Poses as Features. To incorporate poses as tokens into the LLM, it is crucial to align the pose features with a language-contextualized space. To achieve this, we utilize a pretrained Pose-Language model (PoseLM), specifically PoseCLIP, to extract pose features that are aligned with the language domain. The PoseCLIP model comprises a pose backbone [48] and a CLIP text encoder [2], and it undergoes training in two phases. Initially, the pose backbone is pretrained on the NTU RGB+D dataset [49] for action classification. Subsequently, in the second phase, we optimize the similarity between pose features and text features, which encode the prompts describing their action labels, using cross-entropy supervision as outlined in [3]. Further details on the training of this model are provided in Appendix C. These pose features, denoted as $P_{i}\in\mathbb{R}^{F_{p}\times D_{p}}$, where $D_{p}$ represents the pose feature dimension, can be utilized as input tokens for training LLAVIDAL.
3.3 Action-Conditioned Object Cue for LLAVIDAL
To comprehensively understand ADL, it is crucial to not only grasp the semantics of objects but also their trajectories, which are closely linked to the actions performed. Consequently, we propose to explicitly utilize these object trajectories as integral components for training LLAVIDAL. Our framework involves a two-stage pipeline to extract object information directly from RGB video data: (i) Action-conditioned object detection and (ii) Object Localization and Tracking. Both stages leverage off-the-shelf models that are effective without the need for additional training, facilitating integration into LLAVIDAL for ADL analysis.

### v3

# Dominick Reilly1    Rajatsubhra Chakraborty1    Arkaprava Sinha1    Manish Kumar Govind1 Pu Wang1    Francois Bremond3,4    Le Xue2    Srijan Das1

@@ Abstract
Current Large Language Vision Models (LLVMs) trained on web videos perform well in general video understanding but struggle with fine-grained details, complex human-object interactions (HOI), and view-invariant representation learning essential for Activities of Daily Living (ADL). This limitation stems from a lack of specialized ADL video instruction-tuning datasets and insufficient modality integration to capture discriminative action representations. To address this, we propose a semi-automated framework for curating ADL datasets, creating ADL-X, a multiview, multimodal RGBS instruction-tuning dataset. Additionally, we introduce LLAVIDAL, an LLVM integrating videos, 3D skeletons, and HOIs to model ADL’s complex spatiotemporal relationships. For training LLAVIDAL a simple joint alignment of all modalities yields suboptimal results; thus, we propose a Multimodal Progressive (MMPro) training strategy, incorporating modalities in stages following a curriculum. We also establish ADL MCQ and video description benchmarks to assess LLVM performance in ADL tasks. Trained on ADL-X, LLAVIDAL achieves state-of-the-art performance across ADL benchmarks. Code and data will be made publicly available at https://adl-x.github.io/.
Figure 1: Web-video trained Large Language Vision Models (LLVM) struggle to understand the fine-grained details and human-object interactions present in Activities of Daily Living (ADL). We propose LLAVIDAL, an LLVM trained with three modalities on our curated ADL-X dataset. ADL-X is derived from trimmed, multi-view ADL videos and is augmented with skeleton and object modalities.
@@ 1 Introduction
This paper aims explores the potential of Large Language-Vision Language Models (LLVMs) to understand Activities of Daily Living (ADL) which present various challenges including multiple exo-centric viewpoints, fine-grained activities with subtle motion, complex human-object interactions, and long-term temporal relationships. Recently,  [35, 38, 78, 45, 56, 39, 73] have integrated videos into LLMs, leading to the development of video-based LLVMs capable of capturing spatio-temporal features. However, these models are predominantly trained on large-scale web videos [30, 22], which mainly consists of sports clips [8], movie scenes [4, 65], and instructional videos [23, 46, 58]. These videos, typically filmed by professionals, follow strict temporal sequences in closely controlled background (e.g., Paragliding). The evident temporal structure and scene semantics in such videos facilitate spatial understanding within LLVMs, as shown in Figure 1. In contrast, ADL videos pose additional challenges, characterized by lack of a strict temporal structure where diverse actions may unfold concurrently within a single sequence [48]. For instance, a person cooking could intermittently engage in unrelated activities like making a phone call or drinking water, disrupting the linear progression of the composite action cooking. Consequently, existing LLVMs trained on web videos struggle to capture such visually perplexing dynamics inherent in ADL scenarios. Moreover, unlike specialized video architectures designed for understanding ADL [6, 5, 19, 20, 55, 24, 69], these LLVMs lack explicit utilization of cues like 3D skeletons or human-object interactions (HOIs), which are crucial for understanding ADL. These cues facilitate the learning of view-invariant representations and capture fine-grained details essential for interpreting complex human activities. Hence, the current limitations in understanding ADL stem from the lack of instruction tuning of LLVMs on real-world multiview ADL datasets captured in indoor settings and the simplistic design of LLVMs with holistic operations.
To address these challenges, we first propose a semi-automated framework for curating ADL video instruction tuning data. Unlike existing framework [45, 11], we introduce novel strategies to guide AI annotators in focusing on human-localized spatial regions, incorporating temporal unstructuredness in training distribution, and minimizing hallucinations through weak supervision to generate of high-quality video instruction pairs for training of LLVMs. The result of this framework is ADL-X dataset, comprising 100K untrimmed RGB video-instruction pairs, 3D skeletons (S), language descriptions. We then introduce the Large LAnguage VIsion model for Daily Activities of Living (LLAVIDAL), trained on ADL-X. LLAVIDAL integrates synchronized multimodal inputs from videos, including 3D skeletons and HOI (Human-Object Interaction) cues, into the LLM embedding space. Our study explores multiple strategies for multimodal integration in LLAVIDAL, addressing the challenge of jointly aligning all modalities with the LLM embeddings. To tackle this, we propose a Multimodal Progressive (MMPro) training strategy, enabling effective training of LLAVIDAL. Furthermore, we introduce the ADL Multiple Choice Question (ADL MCQ) and ADL video description benchmarks, specifically designed to evaluate LLVM effectiveness in understanding ADL. Empirical results show that LLAVIDAL outperforms other LLVMs, including those trained on datasets ten times larger, on these ADL benchmarks. To summarize our contributions:
@@
  • We introduce ADL-X, the first multiview RGBS instruction tuning dataset for ADL tasks, curated via a novel semi-automated framework for training LLVMs.
  • We present LLAVIDAL, the first LLVM designed for ADL, integrating 3D skeletons and HOI cues into the LLM embedding space, and propose the MMPro training strategy for synchronized multimodal training. LLAVIDAL trained on ADL-X outperforms all baseline LLVMs.
  • We establish new benchmarks, ADL MCQ and video description tasks, to assess ADL understanding in LLVMs.
2 Related Works
Towards emulating human cognitive perception in digital intelligence, initial efforts focused on integrating vision and language modalities [54, 62, 49, 26]. Subsequently, the success of LLMs like GPT [7], PALM [14], BLOOM [72] led to the introduction of multimodal conversational models[81, 66, 40, 74, 1, 12, 2] that combine image pixels and LLMs, we call LLVMs. In this section, we briefly discuss methodologies relevant to developing video conversational models.
Image captioners + LLM. With the emergence of foundation models [52, 76, 34, 70], a natural extension is to use pre-trained VLM encoders to map visual inputs into language representations for LLM processing. For instance, CogVLM [68] uses this approach for image captioning, while Socratic Models [75] and VideoChat [35] adapt it for video captioning. Similarly, dialog-based models like VideoChatCaptioner [10] summarize videos via interactions between ChatGPT [7] and a captioner such as BLIP2 [33], while ChatVideo [67] uses task-specific foundation models with ChatGPT [7] to generate responses to user queries. Efficient video processing approaches, which encode video segments or frames via VLMs followed by single or multiple LLM calls with temporal hierarchy, have been extensively explored in [77, 71, 29, 53, 51]. However, these models rely on static image-to-language mappings, lacking explicit modeling of temporal information.
LLVMs. To align image pixels with the LLM embedding space, image-based methods [81, 74, 17, 40, 2] employ a visual connector, often a simple linear projection layer. In contrast, models like Flamingo [1] and BLIP2 [33] use cross-attention mechanisms, such as a Query Transformer (Q-Former), to align visual features with textual prompts. For video alignment, LLVMs similarly use either linear projectors [45, 11, 28] or Q-Formers [35, 78, 56] for feature encoding and alignment. However, these LLVMs generally capture a holistic view without focusing on fine-grained details or learning view-invariant representations, which are essential for understanding ADLs. In LLAVIDAL, additional video-derived modalities enable its visual connector to address these challenges.
More Modalities + LLVMs. Methods like VideoLLama [78] and VideoLLaVA [38] incorporate additional modalities such as audio and images alongside videos. Similarly, X-InstructBLIP [50], aligns multiple modalities individually with language embedding space, as in LanguageBind [80]. In contrast, we focus on synchronously aligning multiple modalities, requiring joint rather than separate alignment. This is achieved through our MMPro training strategy, which uses progressive training [32, 60] to sequentially integrate modalities into the LLM embedding space.
Figure 2: ADL-X dataset curation pipeline. The ADL-X dataset is derived from the NTU RGB+D 120 dataset through the use of three techniques: Person Augmented Generation, Temporal Stitching, and Weakly Supervised Video Descriptions. The pipeline leverages CogVLM [68] for frame-level caption generation and GPT-3.5 Turbo [7] for summary synthesis and question-answer pair generation.
3 Background: Training LLVM
@@
In this section, we briefly present the general framework for training LLVMs. Given an input video $V_{i}\in\mathbb{R}^{T\times H\times W\times C}$, we encode each frame using the pretrained vision-language model (VLM) CLIP-L/14 [52] to obtain frame-level embeddings $f_{i}\in\mathbb{R}^{T\times h\times w\times D_{v}}$, where $D_{v}$ is the embedding dimension, and $h=H/p$, $w=W/p$ are dimensions adjusted by the patch size $p$. We extract temporal and spatial features by aggregating these embeddings along their respective dimensions [45]. The resulting video-level features, $\mathcal{X}^{v}_{i}\in\mathbb{R}^{F_{v}\times D_{v}}$, where $F_{v}$ denotes the number of spatio-temporal tokens, are obtained by concatenating the aggregated features.
@@
4 ADL Video-instruction Pairs (ADL-X)
This section describes the weakly supervised semi-automated data curation framework employed for the creation of a novel dataset, ADL-X. This curation framework is carefully designed to enable the instruction tuning of LLVMs within the ADL domain. ADL-X comprises video recordings of ADLs from the NTU RGB+D 120 (NTU) dataset [41]. This selection was motivated by the dataset’s focus on ADL videos and its inherent diversity in terms of actions, subjects, and camera viewpoints.
To facilitate LLVM training, question-answer (QA) pairs pairs were systematically generated targeting various aspects of ADL. The semi-automated data curation process for ADL-X incorporates three novel techniques as illustrated in Figure 2: (i) Person Augmented Generation, which enhances video descriptions with spatially focused human related contents; (ii) Temporal Stitching, which connects short action clips for continuity; and (iii) Weakly Supervised Video Descriptions, which facilitates the generation of high-quality descriptions that guide the creation of video QA pairs. These QA pairs address diverse dimensions of ADLs, including human pose configurations, objects related to human actions, scene appearance, and the fine-grained actions performed. Subsequent sections detail the chronological steps undertaken to develop ADL-X.
Person Augmented Generation (PAG). ADL tasks necessitate focusing spatially on the individual’s postures and their interactions with objects, distinct from the contextual background typical in web videos [19]. Specifically, we utilize skeleton data to crop bounding boxes around individuals. This person augmentation strategy minimizes background information in the video frames, effectively diminishes the generation of extraneous content by LLMs in video descriptions, focusing on human actions.
Temporal Stitching (TS). Real-world ADL videos videos typically lack temporal structure, in contrast to instructional videos like cooking, where actions are sequentially linked. To mimic the inherent randomness of ADLs, we develop a methodology to construct long, untrimmed video sequences by stitching together shorter clips. Initially, we generated 160 composite action sequences using GPT prompts to combine individual actions from the NTU dataset’s 120 actions ($A_{1}$, $A_{2}$, …, $A_{120}$). An example of such a sequence is drink water $\rightarrow$ eat snack $\rightarrow$ wipe face. The corresponding short clips from the NTU dataset are then stitched together, ensuring that all clips within a sequence maintain the same subject for consistency. To increase diversity and minimize bias towards specific subject-action pairings, we randomize both the sequence of actions and the subject assignments. This procedure yields 16,206 stitched videos, with a maximum of 7 actions per video.
Weakly Supervised (WS) Video Descriptions. The generation of video descriptions, integral to automating QA pair creation, involves a two-step process: initially, image captions are generated for each frame in the video at a rate of $0.5$ fps using CogVLM [68]. Subsequently, these frame-level captions are synthesized into a cohesive video description via one LLM call. However, this step is prone to introducing hallucinations, potentially degrading the dataset’s quality. To mitigate this, we employ weak supervision by incorporating the action sequence obtained from the short clips into the GPT-3.5 Turbo model, guiding it to generate a structured, coherent description limited to 300 words. This method of weak supervision helps eliminate irrelevant noise from the caption generation process. For details on our prompting strategy, see Appendix G.
Generating QA Pairs. To generate domain-specific QA pairs for ADL, we utilize the dense video descriptions produced in the preceding step. An instruction template (detailed in Appendix G), guides the GPT-3.5 model to formulate questions across various ADL-relevant categories. These categories include video summary, performed actions, spatial details, human-object interactions and other video-specific inquiries. Through this prompting approach, we curated a dataset comprising 100K video instruction pairs, named ADL-X, derived from the stitched ADL videos. Note that this data curation framework, applied to ADL-X, is adaptable and can be extended to other existing datasets for domain-specific training of LLVMs.
5 Multi-modalities of ADL-X
Following the literature for understanding ADL [5, 55], we recognize the importance of integrating additional modalities $\mathcal{M}_{m}$, such as 3D skeletons ($\mathcal{M}_{s}$) and human-object interactions ($\mathcal{M}_{o}$), to enhance the understanding of ADLs. These modalities are critical as ADLs predominantly involve movements of essential body parts or joints, facilitating the learning of view-invariant representations. Moreover, understanding not only the semantics of objects but also their trajectories — integral to the actions performed — is essential for developing fine-grained representations. However, the question remains: how should these modalities be integrated within the LLM of the LLVM?
Figure 3: The multiple modalities of ADL-X. Left: Extraction of skeleton data features ($\mathcal{M}_{s}$) using SkeletonCLIP; Middle: Pipeline for extracting Human-Object Interaction (HOI) features through action-conditioned object detection, localization, and tracking; Right: Outline of our approach for obtaining $\mathcal{M}_{m}$ as QA and $\mathcal{M}_{m}$ as context.
To address this, we investigate several strategies to incorporate skeletons and object tracks into the LLM input space. These include utilizing features extracted from dedicated modality-language encoders, augmenting QA pairs with modality-specific information, and contextualizing inputs to the LLM with modality-specific language as illustrated in Figure 3.
$\mathcal{M}_{m}$ as features. To integrate new modalities into the LLVM input, we map modality-specific data from $\mathcal{M}_{m}\rightarrow\mathcal{X}_{i}^{m}$, which is combined with visual features $\mathcal{X}_{i}^{v}$. It is critical that the features $\mathcal{X}_{i}^{m}$, extracted from skeletons and HOI, are aligned to the language domain to facilitate their integration with the LLM embeddings, analogous to visual features derived from VLM [31]. We provide a detailed methodology for extracting these modality-specific features $\mathcal{X}_{i}^{m}$.
Skeleton Features (SF) - To extract features from the skeleton data $\mathcal{M}_{s}$ to be fed as input to the LLVM, we employ a skeleton-language model, specifically SkeletonCLIP [63]. SkeletonCLIP is a dual-encoder framework that combines a skeleton backbone [79] and a frozen CLIP text encoder [52]. Initially, the skeleton backbone is pretrained on trimmed NTU clips [59] for skeleton action classification. Subsequently, it is fine-tuned to enhance the alignment between skeleton features and language descriptions of actions using cross-entropy supervision. The resulting skeleton features are denoted as $\mathcal{X}^{s}_{i}\in\mathbb{R}^{F_{s}\times D_{s}}$, where $D_{s}$ indicates the dimension of skeleton features. These features are used as input tokens to LLVMs.